A thin film of soap with n = 1.36 hanging in the air reflects dominantly red light with λ = 695 nm.

What is the minimum thickness of the film?
1.278×102 nm

Now this film is on a sheet of glass, with n = 1.54. What is the wavelength of the light in air that will now be predominantly reflected?
Answer #1

Solution)

We know, for constructive interference

Thickness, t=(2m+1)*lambda/4n

T=(2*0+1)*(695*10^-9)/4*1.36= 1.27*10^-7 m=1.278*10^2 nm

======

Here, lambda= 2nt/m =2*1.36*(1.27*10^-7)/1= 3.47*10^-7 m

======÷÷